one face of a cubed-shape package has an area of 7 and five sixth inches.what is the closest total suface area of the package
Found 2 solutions by stanbon, shree840:
Answer by stanbon(75887)

one face of a cubed-shape package has an area of 7 and five sixth sq in.
What is the closest total suface area of the package
-----
There are six faces on a cube.
Total area = 6(7 5/6))
= 6(47/6)
= 47 sq. in
==============
Cheers,
Stan H.

Answer by shree840(260)

Total surface area will be 6times one surface
6*7 5/6
6*47/6=47 Ans